OpenStudy (sugarmochi):

Mary sat for ten tests this term and has an average score of 68. What mark must she gain in the next one to raise her average to 70?

OpenStudy (aabomosalam1998):

Average score= Sum of scores / number of tests taken Average score for ten tests=68=sum of scores/10--> sum of scores=680 By taking another test, the number of tests taken would sum to 11 instead of 10 given the sum of scores in 10 tests and the average score you can easily work out the mark required to gain an average score of 70.

OpenStudy (pythagoras123):

Sum of scores from 10 tests = 680 Let the score required be x, such that x fufils the equation \[\frac{ 680+x }{ 11 }=70\] \[680+x=770\] \[x=90\]